• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 14:40:22 on October 12, 2006

    Entry number 186911

    keyword=Compton slow control IOC rebooted

    IOC for Compton was down. I tried remotely to power cycle it but it did not work. Though logging into the portserver

    telnet cptsv1 2008

    the IOC seemed to be alive.

    Power cycled it again with a "epics reboot" and it came back fine.

    The electron detector was in and is out now.